# FT-Shield: A Watermark Against Unauthorized Fine-tuning in Text-to-Image Diffusion Models

### v1

@@ 3.3 Watermark detector
Augmentation with Generated Images. Given that the size of the protected dataset can be very small, the watermark detector trained on only the protected dataset is unreliable. To address this issue, we propose using generated images from the fine-tuned text-to-image models to augment the training set of the detector. This process unfolds as follows: we first fine-tune the text-to-image diffusion model using a particular fine-tuning method with both clean and watermarked datasets, yielding two separately fine-tuned models. Subsequently, these models are employed to generate two distinct sets of data. Data generated from the model fine-tuned using clean data are incorporated to the original clean dataset, and the images generated from the model fine-tuned with watermarked images are utilized to augment the watermarked dataset for the training of the detector.
If the data protector has the knowledge about the specific fine-tuning method the offender uses, he/she can directly use the images generated from that specific fine-tuning method to augment the data for the classifier training. In the case that the data protector has no knowledge on the fine-tuning method, we have experiments in 4 to demonstrate that with a relatively higher watermark budget, the detector for one fine-tuning method can exhibit a decent adaptability when applied to the data generated by an alternative fine-tuning method. This transferability is notably prevalent in the classifier trained with the augmentation of the data generated by LoRA.
@@ 4 Experiment
In this section, we evaluate the efficacy of our FT-Shield across various fine-tuning methods, subject transfer tasks and different datasets. We first introduce our experimental setups in Section 4.1. In Section 4.2, we evaluate and analyze our approach in terms of its detection accuracy and influence on the image quality during both the protection and audit stages. Then we further investigate our approach from Section 4.3 to Section 4.5 in terms of its performance under fewer fine-tuning steps, transferability across different fine-tuning methods and robustness against image corruptions. And we have demonstrated the ablations studies in  4.6.
4.1 Experiment settings
@@
Model, Task, Dataset and Baselines. We conducted our experiments using the Stable Diffusion as the pretrained text-to-image model. The size of all the images is 512x512. The primary focus was on two tasks: style transfer and object transfer. For style transfer, we utilized 10 datasets sourced from WikiArt, each containing between 20 to 40 images. The object transfer task was more diversified, incorporating datasets of two lifeless objects, and three individual human faces, with each dataset containing five images. We adopted different fine-tuning methods for different tasks: for style transfer, the methods included DreamBooth (Ruiz et al., 2023), Textual-inversion (Gal et al., 2022), Text-to-Image Fine-tuning (von Platen et al., 2022), and LoRA (Hu et al., 2021), while for object transfer, only DreamBooth and Textual-inversion were utilized because the performance of the other two methods is not good. We considered the watermarking methods Inject Memorization (IM) (Wang et al., 2023) and Gen-Watermark (GM) (Ma et al., 2023) which are also proposed for copyright protection against the fine-tuning of text-to-image diffusion model as our baselines.
@@
Implementation Details. For watermark generation, we consider watermark budgets of 4/255 and 2/255 for each dataset. The watermarks are trained with 5-step PGD (Madry et al., 2018) with step size to be 1/10 of the $l_{\infty}$ budget. For watermark detection, we adopted ResNet18 (He et al., 2016) with the Adam optimizer, setting the learning rate to 0.001 and the weight decay to 0.01. In detection stage, we use 60 and 30 prompts for image generations in style transfer and object transfer tasks, respectively. Details about the prompts and hyperparameters of the fine-tuning methods are in Appendix C, and  B.
Evaluation Metrics. We evaluate the effectiveness of our watermarking approach from two perspectives: 1) its influence on the quality of the released protected images and the generated images; 2) its detection accuracy on the generated images of fine-tuned text-to-image models.
@@
  • Image Quality. We use FID (Heusel et al., 2017) to assess the impact of watermarks on image quality. Specifically, we measure the visual discrepancies between the original and watermarked images to evaluate its influence on the released images’ quality. And we also calculate the FID between the images generated from models fine-tuned on clean images with those generated from models fine-tuned on watermarked images to measure the watermark’s influence on the generated images. A lower FID indicates better watermark invisibility.
  • Detection Accuracy. The efficacy of the watermark detection on the generated images is assessed by two metrics. First, the detection rate (or the true positive rate, TPR) quantifies the proportion of instances where the detector accurately identifies images produced by models fine-tuned on watermark-protected images. Second, the false positive rate (FPR) represents the percentage of instances where the detector mistakenly flags images without watermarks as watermarked.
Table 1: Detection accuracy of watermark
ours ($\eta=4/255$) ours ($\eta=2/255$) GW IM
@@
Style DreamBooth 99.50% 0.18% 98.68% 0.87% 93.31% 3.81% 84.27% 4.18%
Textual Inversion 96.12% 3.03% 93.55% 5.25% 78.75% 12.70% 68.07% 0.25%
Text-to-image 98.77% 1.28% 96.77% 3.54% 75.41% 30.03% 71.87% 5.30%
LoRA 97.65% 2.67% 93.37% 6.17% 67.28% 22.72% 67.97% 9.78%
Object DreamBooth 98.93% 1.23% 97.60% 1.13% 91.39% 3.50% 75.97% 1.20%
Textual Inversion 97.73% 1.67% 97.23% 1.97% 88.22% 3.95% 58.41% 22.20%
@@
4.2 Main results
In this section, we show that our FT-Shield can achieve good performance in two aspects: 1) a high detection accuracy; and 2) a small influence on the quality of images. In experiments, we consider the scenario that the specific fine-tuning method of the suspected model is known by the data protector and the protector can use the generated images of that method as augmentation to train the watermark detector. Note that in Section 4.4, we will demonstrate the transferability o our watermark detector that can ensure its effectiveness when the fine-tuning method is unknown to the data protector. In the evaluation of detection accuracy, we measure the TPR and FPR of the watermark detector in the generated images. And we use FID to evaluate the influence of the watermark in the released and generated images. The average of the main metrics across multiple datasets for the two transfer tasks are demonstrated in Table 1 and Table 2, respectively. More statistics about the TPR and FPR are shown in Figure 4. Examples of the watermarked released images and generated images can be found in Figure 3.
@@ 4.2 Main results
According to the results in Table 1 and Table 2, our method is able to protect the images with the highest TPR and lowest FID values in released dataset among all the fine-tuning methods in both style transfer and object transfer tasks. With an $l_{\infty}$ budget to be 4/255, the TPR of the watermark detection can achieve nearly 100% across all the fine-tuning methods. At the same time, the FPR is also very close to 0. Even constrained by a small budget with an $l_{\infty}$ norm to be 2/255, FT-Shield can still achieve a TPR consistently higher than $90\%$ and an FPR no higher than $7\%$. As shown in Figure 4, in general, the variance of the TPR and FPR of our FT-Shield are consistently smaller than the baseline methods, indicating a more stable ability in protecting the images. Although in some cases the FID of the generated images of DreamBooth and Textual Inversion is relatively higher than the baseline methods, our method consistently demonstrates a much higher watermark detection accuracy. This ensures the success of the detection of unauthorized use. In comparison, baseline methods Gen-Watermark (Ma et al., 2023) and IM (Wang et al., 2023) has a decent TPR and FPR in DreamBooth, but both of them cannot achieve a good accuracy in the images generated by other fine-tuning methods. This indicates that they are not able to provide reliable protection against those fine-tuning methods.
@@
FT-Shield ($\eta:4/255$) FT-Shield ($\eta:2/255$) GW IM
Style Released 20.80 6.79 58.04 65.50
@@
Figure 3: Examples of watermarked images (first line) and generated images (other lines) in the style of artist Beihong Xu. The prompt of generation: A frog on a lotus Leaf in the style of [V]
Figure 4: TPR (green/blue boxes) and FPR (orange/red boxes) of watermark detector
@@ 4.3 Performance under insufficient fine-tuning steps
In this subsection, we provide more evidence that the watermarks generated by our FT-Shield can be better assimilated by the fine-tuning process of diffusion models. Based on DreamBooth, we conduct model fine-tuning with much fewer steps compared with standard experiments. And we apply our watermark detector generated in the standard experiment to calculate
@@
Steps Ours (4/255) GW IM
10 56.17% 32.67% 3.50%
20 66.00% 33.50% 7.33%
50 65.96% 52.67% 2.00%
100 66.94% 41.00% 1.83%
200 76.50% 57.97% 13.33%
300 97.17% 66.67% 54.67%
900 100.00% 99.72% 93.83%
the detection rate (TPR) of the watermark. These experiments are mainly done with the style transfer tasks using the paintings of artist Claude Monet. The results are demonstrated in Table 3. As shown in Table 3, FT-Shield consistently achieves the highest detection rate of the watermark when the fine-tuning steps are insufficient. Even when the fine-tuning steps are as few as 10, the detection rate of watermark can achieve 56%. With only 300 steps, the watermark detection can achieve nearly 100%. In comparison, the two baseline methods require much more steps to achieve a high detection rate. This indicates that our FT-Shield can be assimilated by the text-to-image model much earlier than the two baseline methods.
Table 4: Transferability of the watermark detectors
FT Method that the classifier trained on Ours ($4/255$) Ours ($2/255$) GW
DreamBooth 78.51% 75.42% 64.60%
Textual Inversion 84.44% 62.10% 56.52%
Text-to-Image 88.32% 78.48% 68.27%
LoRA 86.42% 82.83% 65.29%
Avg. 84.42% 74.71% 63.67%
4.4 Transferability across different fine-tuning methods
In this subsection, we demonstrate that the detector trained with the images generated by one fine-tuning method can achieve good adaptability to the data generated by another fine-tuning method. In Table 4, we evaluate the detection accuracy (detector’s average accuracy in the watermark class and clean class) of the classifier which is trained with the images from one fine-tuning method on the images generated by the other three fine-tuning methods. The experiments are mainly based on the style transfer task and the performance is compared with Gen-Watermark (Ma et al., 2023), which also requires the generated images in developing the watermark detector. According to the results demonstrated in Table 4, our detectors consistently transfer well to other fine-tuning methods, achieving average accuracy of about 85% and 75% for budgets 4/255 and 2/255, respectively. Notably, the classifier trained with the images from LoRA achieves a transfer accuracy higher than 80% for both budgets. In comparison, the classifiers from GW exhibit inferior adaptability. Each of their classifiers achieves a transfer accuracy lower than 70%.

### v2

@@ 3.3 Mixture of Watermark Detectors
Two-stage training. Instead of training the experts and Gating Model simultaneously, we employ a two-stage training strategy for watermark detection. Specifically, we first train each expert individually with the binary cross-entropy loss. Following the work of Ma et al. (2023), we enhance the training dataset for each expert with images generated by fine-tuned text-to-image models. This process unfolds as follows: we first fine-tune the text-to-image model using a particular fine-tuning method with both clean and watermarked datasets, yielding two separately fine-tuned models. Subsequently, these models are employed to generate two distinct sets of data. Data generated from the model fine-tuned with clean data are incorporated into the original clean dataset, and the images generated from the model fine-tuned with watermarked images are utilized to augment the watermarked dataset for the training of the detector. After the experts are developed, we then train the Gating Model with a cross-entropy loss function, leveraging a dataset organized into various classes, where each class contains images generated by a specific fine-tuning method.
This two-stage training strategy has several advantages. First, it provides good adaptability to new fine-tuning methods for text-to-image models. Updating the MoE system requires only the training of the new experts and the Gating Model. This enables the straightforward integration of previously trained experts without necessitating their retraining. Second, it effectively prevents the potential “collapse problem” (Shazeer et al., 2017) of MoE, which refers to the situation that the prediction relies on only the output of a single expert. Training the experts separately and initially helps to prevent this issue, promoting a more equitable and effective engagement of all experts. Third, it has a good memory efficiency. Since each expert is trained and inferred independently, it is unnecessary to load all experts into memory at once.
@@ 4 Experiment
In this section, we evaluate the effectiveness of FT-Shield across various fine-tuning methods, subject transfer tasks and different datasets. We first introduce our experimental setups in Section 4.1. In Section 4.2 and Section 4.3, we evaluate and analyze the detection performance of FT-Shield. Then we assess the impact of FT-Shield on image quality in Section 4.4. We further investigate our approach in Section 4.5 and Section 4.6 for its performance under fewer fine-tuning steps and robustness against image corruptions. Due to space limitation, our ablation studies on FT-Shield’s performance with a reduced watermark rate and without training the detector on images from fine-tuned models are detailed in Appendix 7.
4.1 Experimental settings
@@
Model, Task and Dataset. We use Stable Diffusion as the pre-trained text-to-image model. The image size is $512\times 512$. We mainly focus on two tasks: style transfer and object transfer. Within style transfer, we explore two subtasks: one centers on art, utilizing 10 diverse datasets from WikiArt, each containing 20 to 40 images; the other focuses on a popular culture, employing the Pokémon BLIP captions dataset (Pinkney, 2022), which includes 833 images. The object transfer task involves two datasets of lifeless objects from Ruiz et al. (2023) and three datasets of individual human faces from CelebA (Liu et al., 2015), each comprising five images. We adopt different fine-tuning methods for different tasks. For style transfer, the methods include DreamBooth (Ruiz et al., 2023), Textual-inversion (Gal et al., 2022), Text-to-Image Fine-Tuning (von Platen et al., 2022), and LoRA (Hu et al., 2021), while for object transfer, only DreamBooth and Textual-inversion are utilized because the performance of the other two methods is not satisfying.
Baselines and Scenarios. Our baselines include Gen-Watermark (Ma et al., 2023) and DIAGNOSIS (Wang et al., 2024a), which are also watermarking methods for IP protection against the fine-tuning of text-to-image model. The two baselines involve different settings in watermark detection. Gen-Watermark (Ma et al., 2023) requires images generated from fine-tuned models to develop watermark detectors, resulting in the requirement of different watermark detectors tailored to different fine-tuning methods. In contrast, DIAGNOSIS (Wang et al., 2024a) utilizes a general watermark detector to evaluate images generated by different fine-tuning methods. It does not require the knowledge about which specific method is used to generate the suspected image.
To compare with the two baseline methods, we evaluate FT-Shield considering two scenarios: 1) the data protector knows the specific fine-tuning method used by the offender; and 2) the protector is unaware of the fine-tuning method. In the first scenario, the protector can directly apply the watermark detector tailored to that fine-tuning method. We denote our method in this setting as FT-Shield-Specific and compare it with Gen-Watermark (Ma et al., 2023). In the second case, we employ the watermark detection based on MoE, denoted it as FT-Shield-MoE, and compare it with DIAGNOSIS (Wang et al., 2024a).
@@
Implementation Details. For watermark generation, we consider watermark budgets of 4/255 and 2/255 for each dataset. The watermarks are trained with 5-step PGD (Madry et al., 2018) with the step size to be 1/10 of the budget. For training the experts and Gating Model used in MoE, we adopt ResNet18 (He et al., 2016) with the Adam optimizer (Kingma & Ba, 2014), taking the learning rate as 0.001 and the weight decay as 0.01. In the detection stage, we use 60 and 30 prompts for image generations in style transfer and object transfer tasks, respectively. Details about the prompts and hyperparameters of the fine-tuning methods are in Appendix E and B.
@@
Evaluation Metrics. We evaluate FT-Shield from two perspectives: 1) its detection performance on data generated by fine-tuned models and 2) its influence on the quality of the released protected images and the generated images. For Detection Performance, we consider two metrics. First, the detection rate (or true positive rate, TPR) quantifies the proportion of instances where the detector accurately identifies images produced by models fine-tuned on watermarked images. Second, the false positive rate (FPR) indicates the rate of instances where the detector mistakenly flags images without watermarks as watermarked. For Image Quality, we use FID (Heusel et al., 2017) for evaluation. Specifically, we measure the visual discrepancies between the original and watermarked images to evaluate its influence on the released images’ quality. We also calculate the FID between the images generated from models fine-tuned on clean images with those generated from models fine-tuned on watermarked images to measure the watermark’s influence on the generated images. A lower FID indicates better invisibility.
Table 1: Detection performance of FT-Shield-Specific
ours ($\eta=4/255$) ours ($\eta=2/255$) Gen-Watermark (Ma et al., 2023)
@@
Style (Arts) DreamBooth 99.50% 0.18% 98.68% 0.87% 93.31% 3.81%
Textual Inversion 96.12% 3.03% 93.55% 5.25% 78.75% 12.70%
Text-to-image 98.77% 1.28% 96.77% 3.54% 75.41% 30.03%
LoRA 97.65% 2.67% 93.37% 6.17% 67.28% 22.72%
Style (Pokemon) DreamBooth 99.5% 0.50% 98.67% 0.33% 95.50% 3.67%
Textual Inversion 96.00% 1.67% 94.67% 4.33% 97.83% 2.67%
Text-to-image 100.00% 0.00% 99.83% 0.00% 98.33% 3.54%
LoRA 99.67% 0.00% 99.83% 0.00% 97.71% 3.54%
Object DreamBooth 98.93% 1.23% 97.60% 1.13% 91.39% 3.50%
Textual Inversion 97.73% 1.67% 97.23% 1.97% 88.22% 3.95%
@@
4.2 Effectiveness of FT-Shield-Specific
@@
In this experiment, we evaluate the performance of FT-Shield-Specific. The average of the TPR and FPR across multiple datasets for different transfer tasks are demonstrated in Table 1. According to the results in Table 1, our method is able to protect the images with the highest TPR and lowest FPR among most of the fine-tuning methods in both style and object transfer tasks. With an $l_{\infty}$ budget of 4/255, the TPR nearly reaches 100% across all fine-tuning methods, while the FPR is close to 0. Even constrained by a small budget (2/255), FT-Shield can still achieve a TPR consistently higher than $90\%$ and an FPR no higher than $7\%$. In comparison, Gen-Watermark (Ma et al., 2023) has decent performance in the Pokemon style transfer and object transfer tasks but fails to achieve good performance in the art style transfer task. This indicates that it cannot consistently provide reliable protection across different applications.
Transferability of Tailored Watermark Detector. We explore the transferability of tailored watermark detectors by assessing their performance on images generated by other fine-tuning methods. The performance of FT-Shield-Specific and Gen-Watermark (Ma et al., 2023) on style transfer tasks is shown in Table 2. The results of object transfer tasks are provided in Appendix 7. From Table 2, it can be observed that generally FT-Shield-Specific outperforms Gen-Watermark (Ma et al., 2023) in terms of the transferability across different fine-tuning methods. Nonetheless, FT-Shield-Specific still experiences an obvious performance drop when the detectors tailored for one fine-tuning method are applied to images generated by other methods. This highlights the need for a method to enhance the adaptability of the detector to different fine-tuning methods. The experiments in Section 4.3 will demonstrate that FT-Shield-MoE can mitigate this problem.
Table 2: Transferability of the tailored watermark detectors in Style Transfer (Arts). Each number indicates the classifier’s detection accuracy (average of true positive rate and true negative rate) when trained on images generated by the column’s fine-tuning method and applied to images from the row’s fine-tuning method.
Method DreamBooth Textual Inversion Text-to-Image LoRA
Ours (2/255) DreamBooth 98.91% 68.94% 79.29% 78.04%
Textual Inversion 70.52% 94.15% 57.42% 58.38%
Text-to-Image 87.27% 57.02% 96.62% 91.15%
LoRA 91.50% 63.79% 93.19% 93.60%
Ours (4/255) DreamBooth 99.66% 87.02% 84.38% 83.50%
Textual Inversion 88.27% 96.55% 68.96% 72.48%
Text-to-Image 94.35% 68.60% 98.75% 93.73%
LoRA 97.71% 73.21% 97.54% 97.49%
Gen-Watermark  (Ma et al., 2023) DreamBooth 95.92% 64.01% 53.87% 64.11%
Textual Inversion 68.01% 83.03% 60.23% 75.39%
Text-to-Image 57.10% 63.94% 72.69% 63.77%
LoRA 70.75% 66.29% 54.39% 72.28%
Figure 4: Examples of watermarked images (first line) and generated images (other lines) in the style of artist Beihong Xu. The prompt of generation: A frog on a lotus Leaf.
4.3 Effectiveness of FT-Shield-MoE
Table 3: Performance of FT-Shield-MoE
ours ($\eta=4/255$) ours ($\eta=2/255$) DIAGNOSIS (Wang et al., 2024a)
[…]
Style (Arts) DreamBooth 99.42% 0.83% 97.79% 1.83% 84.27% 4.18%
Textual Inversion 95.37% 0.58% 95.67% 3.04% 68.07% 0.25%
Text-to-image 99.04% 1.79% 96.83% 3.12% 71.87% 5.30%
LoRA 97.21% 2.87% 91.08% 6.79% 67.97% 9.78%
Style (Pokemon) DreamBooth 99.00% 2.17% 98.67% 2.50% 48.11% 0.67%
Textual Inversion 92.33% 1.67% 91.17% 4.17% 57.05% 4.00%
Text-to-image 99.83% 1.17% 99.83% 0.00% 82.05% 6.67%
LoRA 99.67% 0.67% 99.67% 0.67% 83.56% 7.83%
Object DreamBooth 99.33% 1.08% 98.08% 2.17% 75.97% 1.20%
Textual Inversion 97.00% 2.08% 98.17% 1.67% 58.41% 22.20%
In this subsection, we demonstrate the effectiveness of FT-Shield-MoE. The detection performance of FT-Shield-MoE when applied to images generated by different fine-tuning methods is shown in Table 3. According to the results, FT-Shield-MoE demonstrates outstanding performance across various datasets. It consistently outperforms the baseline method DIAGNOSIS (Wang et al., 2024a) across different fine-tuning methods and different transfer tasks. For most of the fine-tuning methods, the TPR is consistently higher than 90% and the FPR is consistently lower than 5%, which is comparable to the performance of FT-Shield-Specific under the scenario that fine-tuning method is known to the protector.
@@ 4.3 Effectiveness of FT-Shield-MoE
FT-Shield ($\eta:4/255$) FT-Shield ($\eta:2/255$) Gen-Watermark  (Ma et al., 2023) DIAGNOSIS  (Wang et al., 2024a)
Style (Arts) Released 20.80 6.79 58.04 65.50
@@
Style (Pokemon) Released 27.93 10.63 57.90 21.03
  Generated DreamBooth 43.53 38.14 32.22 34.86
Textual Inversion 96.98 45.76 67.24 67.21
Text-to-image 32.52 27.22 47.54 27.73
LoRA 39.53 33.52 49.25 38.82
@@
4.4 Influence on images quality
In this subsection, we assess how FT-Shield affects the quality of both protected and generated images. Given that the impact of the watermark on image quality is irrelevant to the watermark detection setting, we conduct a collective comparison of FT-Shield with the two baseline methods. We demonstrate the average of the FID metric for each transfer task across different datasets in Table 4. According to the results in Table 4, FT-Shield consistently achieves the lowest FID values in the released dataset. For the generated data, in most cases it also leads to a lighter influence on image quality. Although, in some cases, the FID of images generated by DreamBooth and Textual Inversion is relatively higher, as discussed in Section 4.2 and 4.3, FT-Shield consistently achieves higher watermark detection accuracy. This guarantees successful detection of unauthorized usage.
To offer a visual perspective, we also provide examples of the watermarked released images and generated images in Figure 4. More visualizations can be found in Appendix C. These visualizations confirm that FT-Shield’s watermark is nearly imperceptible, maintaining the aesthetic integrity of both protected and generated images across various fine-tuning models.
@@ 4.5 Performance under insufficient fine-tuning steps
 steps  FID Ours (4/255) Gen-Watermark (Ma et al., 2023) DIAGNOSIS  (Wang et al., 2024a)
10 86.05 56.17% 32.67% 3.50%
20 83.90 66.00% 33.50% 7.33%
50 67.32 65.96% 52.67% 2.00%
100 59.42 66.94% 41.00% 1.83%
200 49.84 76.50% 57.97% 13.33%
300 45.93 97.17% 66.67% 54.67%
500 34.66 98.17% 85.67% 79.55%
800 35.25 100.00% 99.72% 93.83%
In this subsection, we provide more evidence that the watermarks of FT-Shield can be better assimilated by the diffusion models at the early stage of the fine-tuning process. Based on DreamBooth, we conduct model fine-tuning with fewer steps compared with standard experiments. Then we apply the watermark detector tailored to DreamBooth to calculate the detection rate (TPR) of the watermark. We also apply FID to evaluate the extent to which the model learns the target style in different steps. The FIDs are derived by comparing images generated at each step against those produced at the completion of fine-tuning. The experiments are mainly done with the style transfer tasks using the paintings of artist Claude Monet. The results are demonstrated in Table 5.
As shown in Table 5, FT-Shield consistently achieves the highest detection rate when the fine-tuning steps are insufficient. Even when the fine-tuning steps are as few as 10, the detection rate can achieve 56%. With only 300 steps, the TPR can achieve nearly 100%. In comparison, the two baseline methods require many more steps to achieve a high detection rate. By observing the change of FID, it can be found that the style has been fully assimilated by the model at the step of 500. At this moment, the TPR of FT-shield is close to 100%, while the other two methods only achieve TPR to be 85.67% and 79.55%. This comparison indicates that our FT-Shield provides more robust copyright protection for data.